OpenStudy (anonymous):

Sid brought $200 to the amusement park . The cost of admission is $16.25, and the cost of a lunch is $6.50. After Sid purchased an admission ticket and a lunch ticket for each person in his group, he had $18.25 left. How many people are Sid's group? For my answer i Got that theres 7 people in his group Can someone check to see if im correct/accurate

OpenStudy (ccswims):

i do not think it's 7 cause when i multiplied and added everything up it wasn't 7

OpenStudy (taramgrant0543664):

I'll let x represent the number of people 6.50x+16.25x=200-18.25 22.75x=181.75 x=181.75/22.75 X=8 I think I did the math right

OpenStudy (anonymous):

For the answer choices 8 isn't one of them @taramgrant0543664

OpenStudy (usukidoll):

given: Sid brought $200 to the park and had only $18.25 left so $181.75 was spent price of admission is $16.25 for a sack lunch it's $6.50 so $16.25+$6.50 = $22.75. per person hmmmmmm.

OpenStudy (usukidoll):

181.75/22.75 = 7.98901 the problem is that there is no such thing as a .9 person. -_-! round that number up and it's approx. 8 ???

OpenStudy (usukidoll):

what about... money spent/ number of people ? 181.75 / 4 = 45.44 181.75 / 5 = 36.35 181.75 / 6 = 30.29 181.75 / 7 = 25.96

OpenStudy (anonymous):

To be honest I think the answer is 7

OpenStudy (usukidoll):

181.75/7.9 = 23.00 181.75/8 = 22.71875 = 22.72 if there were 7 people, that guy will still have some spare change $25.96 - 22.75 = $3.21

OpenStudy (usukidoll):

maybe it's one of those common core guestimating problems.

OpenStudy (anonymous):

I really don't know there's probably a error in this problem

OpenStudy (anonymous):

Let's go to a different problem

OpenStudy (usukidoll):

I've read that common core problem have guestimates as answers as opposed to the real thing like for example we all know that 3 x 5 = 15 but according to common core 3 x 4 is 13 or something like that what the heck?!

OpenStudy (usukidoll):

x = 181.75/22.75 = 7.98901 181.75/7.98901 = 22.75000 = 22.75 but a .98 person doesn't make sense. so .. 7 will have to be the closest if we had to respect the answer choices. a la carte common core